# Ballistic electron transport described by a fourth order Schrödinger equation

### v1

@@ Abstract
A fourth order Schrödinger equation for the description of charge transport in semiconductors in the ballistic regime is proposed with the inclusion of non parabolic effects in the dispersion relation in order to go beyond the simple effective mass approximation. Similarly to the standard (second order) Schrödinger equation, the problem is reduced to a finite spatial domain with appropriate transparent boundary conditions to simulate charge transport in a quantum coupler [1, 2, 3], where an active region representing an electron device is coupled to leads which take the role of reservoirs. Some analytical properties are investigated and a generalized formula for the current is obtained. Numerical results show the main features of the solutions of the new model. In particular, an effect of interference appears due to a richer wave structure than that arising for the Schrödinger equation in the effective mass approximation.
@@
Keywords: Fourth Order Schrödinger Equation, Resonant Tunneling Diode, Transparent Boundary Conditions
@@ 1 Introduction
The enhanced miniaturization of modern electron devices makes mandatory to adopt a full quantum description of the charge transport. Among the main approaches for nanoscale devices, in the literature we find the use of the Wigner equation [4, 5, 6, 7, 8], the nonequilibrium Green function [9] and the Schrödinger equation [2, 3]. The latter has been employed for example for the simulation of resonant tunneling diodes [1, 10, 11, 12] in the ballistic regime, giving good characteristic curves, at least from a qualitative point of view. However, the Schrödinger equation is almost always employed in the effective mass approximation [13] obtained with a renomalization of the bare electron mass by a factor leading to a reduced mass that encompass the presence of the periodic potential the charge carriers undergo inside the crystal lattice. The main aim of this article is to include non parabolic effects in the description of electron transport in nanoscale devices in order to go beyond the standard parabolic band approximation within the context of the Schrödinger equation. In fact, already in the semiclassical case the simple parabolic band approximation leads to an overestimation of the current and it is considered as not very accurate [14, 15, 16]. We expect a similar effect also in a full quantum setting. The complete expression of the dispersion relation in a semiconductor can be obtained by numerical approaches [7, 15]. However, some analytical approximations are often adopted with results that improve those based on the simple parabolic band. In this paper, the Kane dispersion relation will be considered. Including the complete Kane dispersion relation in the Schrödinger equation is a daunting task; here we include some first order effects beyond the effective mass approximation obtaining a fourth order Schrödinger equation. Moreover, to take into account the additional non parabolic term in the dispersion relation, we have formulated a generalization of the Transparent Boundary Conditions (TBCs), already devised in [1, 2, 3] for the standard (second order) Schrödinger equation. We observe that in literature fourth order Schrödinger equation have already been considered in other contexts, e.g. in [17] for nonlinear Schrödinger-type equations with higher-order dispersion.
@@
In the present article a one dimensional case will be investigated. We will be able to reduce the problem of solving the Schrödinger equation only in the active area with the new TBCs, obtaining a generalized Sturm-Liouville problem of fourth order. Some analytical properties of the new model are examined. It is proved that the resulting boundary value problem is well posed and an efficient numerical strategy is devised. Numerical results for a single particle under several kinds of potential - single step, single and double barrier, double barrier plus a linear potential - show that the new model reveals new features, in particular an effect of interference due to a richer wave structure than that arising for the Schrödinger equation in the effective mass approximation.
@@
The plan of the paper is as follows. In Sec. 2 the fourth order Schrödinger equation is obtained by including the effect of non parabolicity up to the first order. In Sec. 3 the appropriate open boundary conditions are deduced and in Sec. 4 the well posedness of the resulting boundary value problem is proved under suitable conditions. Sec. 5 is devoted to the expression of the current for the fourth order Schrödinger equation while in the last section several numerical simulations are presented to show the difference with the standard Schrödinger equation in the effective mass approximation.
2 Fourth order Schödinger equation
@@
The general form of the dispersion relation in a semiconductor is obtained by solving the single electron Schödinger equation under a periodic potential by employing the Bloch’s theorem [7, 8, 18]. The complete form can be obtained only numerically, e.g. with the use of pseudo-potential [19, 20]. However, often in the applications analytical approximations are adopted [14]. Among these, one commonly used is the Kane dispersion relation, which in its isotropic form is given in implicit form as
@@
where $\epsilon(k)$ is the electron energy, $k$ is the modulus of the electron wave-vector and $m^{*}$ is the effective electron mass, for example for GaAs $m^{*}=0.067m_{e}$ with $m_{e}$ bare electron mass. The positive parameter $\alpha$ is named non parabolicity factor, since in the limit $\alpha\to 0^{+}$ one recovers the standard parabolic band approximation
@@
The direct use of relation (2.3) into the Schrödinger equation (SE) for electron transport is not an easy task. The most straight-forward approach consists in writing the SE in the momentum representation but in such a case the analogous formulation of the open boundary conditions similar to those in the coordinate representation, see for example [21], is still an open problem. However, since the parabolic band approximation gives results which are not very accurate in the semiclassical case [16, 22], it is desirable to include, in any case, effects due to non parabolicity in the charge transport. To this aim we proceed with a perturbative approach. For semiconductors one has $10<\alpha<$ 1 (in unit of eV-1), so we can expand the dispersion relation as powers of $\alpha$. Formally we consider $\alpha$ as a small parameter. Up to the first order, we get
@@
where $\Psi$ is the stationary wave function, $H=\epsilon(P)-qV(X)$ is the Hamiltonian and $E$ is the eigen-energy, with $V({\bf x})$ the electrostatic potential and $q$ elementary (positive) electron charge. By inserting the approximation of $\epsilon$ up to first order in $\alpha$, we get the following fourth order stationary Schödinger Equation (SE4) in the coordinate representation
@@
and assume the electrostatic potential $V({\bf x})$ as an external (assigned) field.
@@
$\epsilon(k)=\frac{p^{2}}{2m^{*}}-\alpha\frac{p^{4}}{4(m^{*})^{2}}$
vanishes for $k=0$ and for $k=\pm\displaystyle{\sqrt{\frac{2m^{*}}{\alpha}}}$. Therefore, we restrict the momentum to the set
@@
3 Transparent boundary conditions for the fourth order Schrödinger equation
A typical structure of electron device comprises several parts. A schematisation used in several papers [1, 2, 12] includes the semiconductor part plus the contacts which are considered as infinite waveguides, i.e. the access zones. In a one dimensional geometry, which is appropriate for example for a resonant tunneling diode, the problem is posed on the real axis which is considered as the union of the half-line ],0[]$\infty$,0[] - ∞ , 0 [ representing the left contact (region $I$), the active region $[0,L]$ (region $II$) representing the semiconductor device, namely, the area where all the relevant physical phenomena occur, and the half-line $]L,+\infty[$ (region $III$) representing the right contact (see Fig. 1). The contacts are metallic and along them the electrostatic potential is constant. We assume that $V(x)$ is continuous at each interface contact/semiconductor
@@ 3 Transparent boundary conditions for the fourth order Schrödinger equation
and belonging to $L^{\infty}(\mathbb{R})$. In the sequel we set $V(0)=V_{0}=0$ and $V(L)=V_{L}\geq 0$.
@@
We split the solution of SE4 as follows
@@
The TBCs have been obtained for the second order SE (hereafter SE2) for the first time in [1] and several analytical investigations have been performed in [2, 3, 23, 24]. Here we want to generalize such TBCs to the case of SE4. In region $I$ SE4 admits four plane waves, two of them with positive momentum and the other two with negative momentum. Therefore a first difference between SE2 and SE4 is that the reflected part of the wave function is a combination of two plane waves instead of one and as a consequence, two reflection coefficients $r_{1}$ and $r_{2}$ appear. Similarly, in region $III$ we have two transmitting coefficients $t_{1}$ and $t_{2}$.
@@
Indeed, let us consider the one-dimensional fourth order stationary Schrödinger equation
@@
and, under the considered hypotheses, by taking the limit as $\delta\to 0^{+}$ the continuity of $\Psi^{{}^{\prime\prime\prime}}$ at $x=0$ follows. Similar results holds at $x=L$.
@@
If we substitute (3.8), (3.9) in (3.6), (3.7) one obtain the boundary conditions at $x=0$
@@
Here for $\beta\in\mathbb{R}$, $\sqrt{\beta}$ must be intended as the positive square root if $\beta\geq 0$. If we get complex roots ±$\pm(\delta+i\sigma)$ then we have to choose the sign that leads to an evanescent wave in considered region. Therefore, for $x>L$, we have to take the sign $+$.
@@
4 Well posedness of the fourth order Schrödinger equation with open boundary conditions
We want to establish some conditions for the well posedness of SE4 with the open boundary conditions devised in the previous section. Due to the symmetry of the problem we will investigate only the the case $p>0$ but the results are valid also when $p<0$. Existence and uniqueness results have been established in [2, 3] for the classical (second order) Schrödinger equation but SE4 is much more complex to tackle, so at the present time we are able to get the same properties only under suitable conditions. These will be checked numerically in the last section.
In the sequel we will denote with $W^{4,1}(0,L)$ the Sobolev space
@@ 4 Well posedness of the fourth order Schrödinger equation with open boundary conditions
If $k_{1}=k_{2}$ and $k_{3}\neq k_{4}$ the boundary conditions give rise to the following linear system
@@
If $k_{1}\neq k_{2}$ and $k_{3}=k_{4}$ the boundary conditions give rise to the following liner system
@@ 4.1 Example: step function potential
Observe that a step potential naturally appears in hetero-junctions made putting together two semiconductor materials possessing different work functions.
@@
Similarly, for SE4 wee look for fundamental solutions $\varphi_{j},\ j=0,1,2,3$ of the type
@@
The most evident difference is the fact that in the region $[\frac{L}{2},L]$ SE2 predicts a plane wave, which has a constant modulus, while the solution of SE4 is the superposition of two plane waves which produces interference, as well evident from Fig. 2. The difference is more marked in the case $V_{L}=$ 0.3 V. Instead, the solutions to SE2 and SE4 are in good agreement in the interval $[0,\frac{L}{2}]$, although there are some slight discrepancy in the extrema.
Even simple, this example clearly reveals that the non parabolic corrections to the dispersion relation can produce interesting effects such as the observed interference. The passage through the device transforms an incoming electron in a coherent state described by a simple wave to an outgoing scattering state which is the superposition of two plane waves.
@@ 5 Expression of the probability current
The evolution equation for the square modulus of the wavefunction can be written in a divergence form with a flux which represents for unit charge the current of the single electron. We want to derive the expression of the current for the SE4.
@@
Proof. Multiplying (5.1) by $\overline{\Psi}$ we have
@@
In the one dimensional case the current reads
@@
As example, let us consider a plane wave $\Psi=e^{ikx}$ with $k=\frac{p}{\hbar}$. In such a case $J$ can be explicitly written as a sum of two components $J_{0},J_{1}$ with $J_{0}=\frac{\hbar k}{m}$ and $J_{1}=-\alpha\frac{\hbar^{3}k^{3}}{m^{*2}}$. In Fig. 3, the values of $J$ versus the wave-vector is plotted setting $\alpha=0.242$ eV-1. We see that the current is lower than that one has in the effective mass approximation, even if the difference is significant for values of $k$ greater than about 0.4 nm-1.
@@ 6.3 A resonant tunneling diode type potential
A dispersion relation beyond the effective mass approximation has been included in the Schrödinger equation for a single electron. This leads to a hierarchy of Schrödinger equations of increase order. A detailed analysis has been performed in the case of the fourth order Schrödinger equation for describing charge transport in a semiconductor device. An explicit relation of the probability current density has been obtained and appropriate transparent boundary conditions have been devised to reduce the problem to a generalized Sturm-Liouville problem in a finite spacial domain. Conditions for the well-posedness of the resulting boundary value problem have been formulated. The shown examples of solutions highlight that the fourth order Schrödinger equation gives rise to interesting effects of resonance which are missing from SE2. As future application, SE4 will be employed for the simulation of a resonant tunneling diode.

### v2

# Ballistic electron transport described by a fourth-order Schrödinger equation

@@ Abstract
A fourth-order Schrödinger equation for the description of charge transport in semiconductors in the ballistic regime is proposed with the inclusion of non-parabolic effects in the dispersion relation in order to go beyond the simple effective mass approximation. Similarly to the standard (second order) Schrödinger equation, the problem is reduced to a finite spatial domain with appropriate transparent boundary conditions to simulate charge transport in a quantum coupler [9, 7, 8], where an active region representing an electron device is coupled to leads which take the role of reservoirs. Some analytical properties are investigated, and a generalized formula for the current is obtained. Numerical results show the main features of the solutions of the new model. In particular, an effect of interference appears due to a richer wave structure than that arising for the second-order Schrödinger equation in the effective mass approximation.
@@
Keywords: Fourth-Order Schrödinger Equation, Resonant Tunneling Diode, Transparent Boundary Conditions
@@ 1 Introduction
The enhanced miniaturization of modern electron devices makes mandatory to adopt a full quantum description of the charge transport. Among the main approaches for nanoscale devices, in the literature we find the use of the Wigner equation [1, 2, 3, 4, 5], the nonequilibrium Green function [6] and the Schrödinger equation [7, 8]. The latter has been employed, for example, for the simulation of resonant tunneling diodes [9, 10, 11, 12] in the ballistic regime, giving good characteristic curves, at least from a qualitative point of view. However, the Schrödinger equation is almost always employed in the effective mass approximation [13] obtained with a renormalization of the bare electron mass by a factor leading to a reduced mass that encompasses the presence of the periodic potential the charge carriers undergo inside the crystal lattice. The main aim of this article is to include non-parabolic effects in the description of electron transport in nanoscale devices in order to go beyond the standard parabolic band approximation within the context of the Schrödinger equation. In fact, already in the semiclassical case the simple parabolic band approximation leads to an overestimation of the current, and it is considered as not very accurate [14, 15, 16]. We expect a similar effect also in a full quantum setting. The complete expression of the dispersion relation in a semiconductor can be obtained by numerical approaches [4, 15]. However, some analytical approximations are often adopted with results that improve those based on the simple parabolic band. In this paper, the Kane dispersion relation is considered. Including the complete Kane dispersion relation in the Schrödinger equation is a daunting task; here we retain some first order effects beyond the effective mass approximation obtaining a fourth-order Schrödinger equation. Moreover, to take into account the additional non-parabolic term in the dispersion relation, we have formulated a generalization of the Transparent Boundary Conditions (TBCs), already devised in [9, 7, 8] for the standard (second order) Schrödinger equation.
We highlight that in literature fourth-order Schrödinger equations have already been considered in other contexts, e.g., in [17] for higher-order dispersion and in [18] for spin current. Recently some attempts have been also carried on to get a higher-order Schrödinger equation in terms of a real-valued wave function starting from some ideas present in the original paper by Schrödinger himself [19]; with this spirit in [20] a fourth-order equation for a real ”wave function” is proposed, which resembles some analogy with certain equations of elasticity. However, we remark that the underlying idea of these papers is different from the subject of our work because we want to incorporate general dispersion in the canonical Schrödinger equation with a standard (complex-valued) wave function as common in solid state physics.
@@
In the present article a one-dimensional case will be investigated. We will be able to reduce the problem of solving the Schrödinger equation only in the active area with the new TBCs, obtaining a generalized Sturm-Liouville problem of fourth-order. Some analytical properties of the new model are examined. It is proved that the resulting boundary value problem is well posed and an efficient numerical strategy is devised. Numerical results for a single particle under several kinds of potential - single step, single and double barrier, double barrier plus a linear potential - show that the new model reveals new features, in particular an effect of interference due to a richer wave structure than that arising for the Schrödinger equation in the effective mass approximation.
@@
The plan of the paper is as follows: In Sec. 2 the fourth-order Schrödinger equation is obtained by including the effect of non-parabolicity up to the first order. In Sec. 3, the appropriate open boundary conditions are deduced, and in Sec. 4 the well-posedness of the resulting boundary value problem is proved under suitable conditions. Section 5 is devoted to the expression of the current for the fourth-order Schrödinger equation, while in the last section several numerical simulations are presented to show the difference with the standard Schrödinger equation in the effective mass approximation.
2 Fourth-order Schödinger equation
@@
The general form of the dispersion relation in a semiconductor is obtained by solving the single-electron Schödinger equation under a periodic potential by employing the Bloch’s theorem [4, 5, 21]. The complete form can be obtained only numerically, e.g., with the use of pseudo-potential [22, 23]. However, often in the applications analytical approximations are adopted [14]. Among these, one commonly used is the Kane dispersion relation, which in its isotropic version is given in implicit form as
@@
where $\epsilon(k)$ is the electron energy, $k$ is the modulus of the electron wave-vector and $m^{*}$ is the effective electron mass; for example, for GaAs $m^{*}=0.067m_{e}$ with $m_{e}$ bare electron mass. The positive parameter $\alpha$ is named non-parabolicity factor, since in the limit $\alpha\to 0^{+}$ one recovers the standard parabolic band approximation
@@
The direct use of relation (2.3) into the Schrödinger equation (SE) for electron transport is not an easy task. The most straightforward approach consists of writing the SE in the momentum representation but in such a case the analogous formulation of the open boundary conditions similar to those in the coordinate representation, see for example [24], is still an open problem. In any case, since the parabolic band approximation gives results, which are not very accurate in the semiclassical case [16, 25], it is desirable to include effects due to non-parabolicity in the charge transport. To this aim, we proceed with a perturbative approach. For semiconductors, one has $10<\alpha<$ 1 (in unit of eV-1), so we can expand the dispersion relation as powers of $\alpha$. Formally we consider $\alpha$ as a small parameter. Up to the first order, we get
@@
where $\Psi$ is the stationary wave function, $H=\epsilon(P)-qV(X)$ is the Hamiltonian and $E$ is the eigen-energy, with $V({\bf x})$ the electrostatic potential and $q$ elementary (positive) electron charge. By inserting the approximation of $\epsilon$ up to first order in $\alpha$, we get the following fourth-order stationary Schödinger Equation (SE4) in the coordinate representation:
@@
and assume the electrostatic potential $V({\bf x})$ as an external (assigned) real field.
@@
$\epsilon(p)=\frac{p^{2}}{2m^{*}}-\alpha\frac{p^{4}}{4(m^{*})^{2}}$
vanishes for $p=0$ and for $p=\pm\displaystyle{\sqrt{\frac{2m^{*}}{\alpha}}}$. Therefore, we restrict the momentum to the set
@@
It is interesting to remark that the fourth-order SE derived in [20] following an approach devised by Schrödinger himself differs from (2.5) for several aspects. The equation proposed in [20] contains also first and second derivatives of the potential but, more stringently, has solution which are real-valued. We have only included a more general dispersion relation in the framework of the canonical quantum wave equation, so the (more precisely pseudo-differential) operator related to the electron energy band acts only on the wave function and no derivatives of the potential arise.
3 Transparent boundary conditions for the fourth-order Schrödinger equation
A typical structure of an electron device comprises several parts. A schematization used in several papers [9, 7, 12] includes the semiconductor part plus the contacts, which are considered as infinite waveguides, the access zones. In a one-dimensional geometry, which is appropriate, for example, for a resonant tunneling diode, the problem is posed on the real axis which is considered as the union of the half-line ],0[]$\infty$,0[] - ∞ , 0 [ representing the left contact (region $I$), the active region $[0,L]$ (region $II$) representing the semiconductor device, namely the area where all the relevant physical phenomena occur, and the half-line $]L,+\infty[$ (region $III$) representing the right contact (see Fig. 1). The contacts are metallic, and along them the electrostatic potential is constant. We assume that $V(x)$ is continuous at each interface contact/semiconductor
@@ 3 Transparent boundary conditions for the fourth-order Schrödinger equation
and belonging to $L^{\infty}(\mathbb{R})$. In the sequel, we set $V(0)=V_{0}=0$ and $V(L)=V_{L}\geq 0$.
@@
We split the solution of SE4 as follows:
@@
The TBCs have been obtained for the second-order SE (hereafter SE2) for the first time in [9] and several analytical investigations have been performed in [7, 8, 26, 27]. Here we want to generalize such TBCs to the case of SE4. In region $I$ SE4 admits four plane waves, two of them with positive momentum and the other two with negative momentum. Therefore, a first difference between SE2 and SE4 is that the reflected part of the wave function is a combination of two plane waves instead of one and as a consequence, two reflection coefficients $r_{1}$ and $r_{2}$ appear. Similarly, in region $III$ we have two transmitting coefficients $t_{1}$ and $t_{2}$.
@@
Indeed, let us consider the one-dimensional fourth-order stationary Schrödinger equation
@@
and, under the considered hypotheses, by taking the limit as $\delta\to 0^{+}$ the continuity of $\Psi^{{}^{\prime\prime\prime}}$ at $x=0$ follows. Similar results hold at $x=L$.
@@
If we substitute (3.8), (3.9) in (3.6), (3.7) one obtains the boundary conditions at $x=0$
@@
Here for $\beta\in\mathbb{R}$, $\sqrt{\beta}$ must be intended as the positive square root if $\beta\geq 0$. If we get complex roots ±$\pm(\delta+i\sigma)$ then we have to choose the sign that leads to an evanescent wave in the considered region. Therefore, for $x>L$, we have to take the sign $+$.
@@
4 Well-posedness of the fourth-order Schrödinger equation with open boundary conditions
We want to establish some conditions for the well-posedness of SE4 with the open boundary conditions devised in the previous section. Due to the symmetry of the problem, we will investigate only the the case $p>0$ but the results are valid also when $p<0$. Existence and uniqueness results have been established in [7, 8] for the classical (second-order) Schrödinger equation but SE4 is much more complex to tackle, so at the present time we are able to get the same properties only under suitable conditions. These will be checked numerically in the last section.
In the sequel, we will denote with $W^{4,1}(0,L)$ the Sobolev space
@@ 4 Well-posedness of the fourth-order Schrödinger equation with open boundary conditions
If $k_{1}=k_{2}$ and $k_{3}\neq k_{4}$ the boundary conditions give rise to the following linear system:
@@
If $k_{1}\neq k_{2}$ and $k_{3}=k_{4}$ the boundary conditions give rise to the following linear system:
@@ 4.1 Example: step function potential
Observe that a step potential naturally appears in hetero-junctions made by putting together two semiconductor materials possessing different work functions.
@@
Similarly, for SE4 we look for fundamental solutions $\varphi_{j},\ j=0,1,2,3$ of the type
@@
The most evident difference is the fact that in the region $[\frac{L}{2},L]$ SE2 predicts a plane wave, which has a constant modulus, while the solution of SE4 is the superposition of two plane waves which produces interference, as well evident from Fig. 2. The difference is more marked in the case $V_{L}=$ 0.3 V. Instead, the solutions to SE2 and SE4 are in good agreement in the interval $[0,\frac{L}{2}]$, although there are some slight discrepancies in the extrema.
Even though simple, this example clearly reveals that the non-parabolic corrections to the dispersion relation can produce interesting effects such as the observed interference. The passage through the device transforms an incoming electron in a coherent state described by a simple wave to an outgoing scattering state which is the superposition of two plane waves.
@@ 5 Expression of the probability current
The evolution equation for the square modulus of the wavefunction can be written in a divergence form with a flux which represents for unit charge the current of the single electron. We want to derive the expression of the current for the SE4 in the spirit of the Ehrenfest’s theorem.
@@
Proof. Multiplying (5.1) by $\overline{\Psi}$ one has
@@
In the one-dimensional case the current reads
@@
As example, let us consider a plane wave $\Psi=e^{ikx}$ with $k=\frac{p}{\hbar}$. In such a case $J$ can be explicitly written as a sum of two components $J_{0},J_{1}$ with $J_{0}=\frac{\hbar k}{m}$ and $J_{1}=-\alpha\frac{\hbar^{3}k^{3}}{m^{*2}}$. In Fig. 3, the values of $J$ versus the wave-vector are plotted setting $\alpha=0.242$ eV-1. We see that the current is lower than that one has in the effective mass approximation, even if the difference is significant for values of $k$ greater than about 0.4 nm-1.
@@
We remark that a generalization of the expression of the current which takes account the transport of spin with a Pauli formalism can be found in [18].
@@ 6.3 A resonant tunneling diode type potential
A dispersion relation beyond the effective mass approximation has been included in the Schrödinger equation for a single electron. This leads to a hierarchy of Schrödinger equations of increasing order. A detailed analysis has been performed in the case of the fourth-order Schrödinger equation for describing charge transport in a semiconductor device. An explicit relation of the probability current density has been obtained and appropriate transparent boundary conditions have been devised to reduce the problem to a generalized Sturm-Liouville problem in a finite spatial domain. Conditions for the well-posedness of the resulting boundary value problem have been formulated. The shown examples of solutions highlight that the fourth-order Schrödinger equation gives rise to interesting effects of resonance which are missing from SE2. As future application, SE4 will be employed for the simulation of a resonant tunneling diode.